# Conformal prediction for Deep Classifier via Label Ranking

### v1

@@ 3.2 Method
In the analysis above, we demonstrate that removing the probability value in APS can largely decrease the size of prediction sets. On the other hand, the expected value of the set size (shown in Eq. 6) will oscillate between $k-1$ and $k$, after removing the probability value. This implies a shortcoming of the modified non-conformity score in adaptation to instance-wise uncertainty, which may cause overcovering on easy examples.
@@
To alleviate this limitation, we propose a novel conformal prediction algorithm, named Sorted Adaptive Prediction Sets. The key idea behind this algorithm is to minimize the dependence of the non-conformity score on the probability values while retaining the uncertainty information. In particular, we discard all the probability values except for the maximum softmax probability, which is usually used to measure the model confidence in the prediction. Formally, the non-conformity score can be calculated as
@@
$S(\bm{x},y,u;\hat{\pi}):=\left\{\begin{array}[]{ll}u\cdot\hat{\pi}_{max}(\bm{x}),\qquad\text{if\quad}o(y,\hat{\pi}(\bm{x}))=1,\\ \hat{\pi}_{max}(\bm{x})+(o(y,\hat{\pi}(\bm{x}))-2+u)\cdot\lambda,\qquad\text{otherwise},\end{array}\right.$ (8)
@@
where $\lambda$ is a hyperparameter showing the weight of ranking information, $\hat{\pi}_{max}(\bm{x})$ denotes the maximum softmax probability and $u$ denotes a uniform random variable. We provide a detailed analysis on the effect of $\lambda$ in Section 5.
[…]
Figure 2: Distribution of non-conformity scores for examples with wrong predictions.
In Eq. 8, we incorporate the uncertainty information via the maximum probability $\hat{\pi}_{max}(\bm{x})$, and use the constant $\lambda$ to mitigate the undesirable influence of tail probabilities. In this manner, the SAPS method can not only produce sets of small size, but also communicate instance-wise uncertainty. In other words, the prediction set can be smaller for easy inputs than for hard ones. We illustrate with an experiment in Figure 2, where the examples with wrong predictions have higher non-conformity scores provided by SAPS, compared to those of APS and RAPS. Moreover, for examples, with correct predictions, the non-conformity scores defined in APS, RAPS, and APS are equivalent as the ranks of ground-truth labels are 1 (i.e. $S(\bm{x},y,u;\hat{\pi})=u\cdot\hat{\pi}_{max}(\bm{x}))$. The results indicate that the non-conformity score of SAPS can better characterize the deviation between a given example and the training data.
In what follows, we provide a formal analysis to show the effectiveness of our SAPS algorithm. We start from showing the finite-sample marginal coverage properties:
Proposition 1.
(Coverage guarantee of SAPS). Suppose $\left(\bm{x}_{i},y_{i},u_{i}\right)_{i=1,\ldots,n}$ and $\left(\bm{x}_{n+1},y_{n+1},u_{n+1}\right)$ are i.i.d. and let $\mathcal{C}_{1-\alpha}(\bm{x},u)$ be defined as in Eq. 3. Then for a confidence level $\alpha$, we have the following coverage guarantee:
$\mathbb{P}\left(y_{n+1}\in\mathcal{C}\left(\bm{x}_{n+1},u_{n+1}\right)\right)\geq 1-\alpha$
The corresponding proof is provided in Appendix B. In the following, we further prove that SAPS always dominates APS in the size of prediction sets.
Proposition 2.
(SAPS dominates APS) If $\hat{\pi}$ is well-calibrated and $\lambda\geq 1-\frac{1}{K}$, for any test instance $\bm{x}\sim\mathcal{P}_{\mathcal{X}}$ with a significance level $\alpha$,
$\mathbb{E}_{u\sim[0,1]}\{|\mathcal{C}(\bm{x},u)|\}\leq\mathbb{E}_{u\sim[0,1]}\{|\tilde{\mathcal{C}}(\bm{x},u)|\},$
where $u\sim U[0,1]$. $\mathcal{C}(\cdot)$ and $\tilde{\mathcal{C}}(\cdot)$ represent the prediction set from SAPS and APS, respectively.
In other words, SAPS consistently generates a smaller prediction set than APS, while both algorithms maintain the desired marginal coverage rate. In Section 5, we provide an analysis to show the connection between calibration and the performance of SAPS.
4 Experiments
4.1 Experimental Setup
Classification datasets. We consider three prominent datasets in our study: ImageNet (Deng et al., 2009), CIFAR-100 and CIFAR-10 (Krizhevsky et al., 2009), which are common benchmarks for conformal prediction. In the case of ImageNet, we split the test dataset containing 50000 images into 30000 images for the calibration set and 20000 images for the test set. For CIFAR-100 and CIFAR-10, We divide the corresponding test dataset equally into a calibration set containing 5000 images and a test set containing 5000 images.
Models. We employ twelve different classifiers, including nine standard classifiers, two transformer-based models, i.e., ViT (Dosovitskiy et al., 2020) and DeiT (Touvron et al., 2021), and a Vision-Language Model named CLIP (Radford et al., 2021). Aside from CLIP with zero-shot prediction capabilities, the remaining models are the pre-trained models on ImageNet. For CIFAR-10 and CIFAR-100, these models will be fine-tuned on the pre-trained models. Moreover, all classifiers are calibrated by the Temperature scaling procedure (Guo et al., 2017).
Conformal prediction algorithms. We compare the proposed method against APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021b). We select the hyper-parameters for RAPS from a set, i.e., $\{0.001,0.01,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$. For SAPS, the parameters set of $\lambda$ is an arithmetic sequence between 0.02 and 0.6, spaced 0.03 apart. We choose the hyper-parameter that achieves the smallest set size on a validation set, which is a subset of the calibration set. All experiments are conducted with ten trials, and the median results are reported.
Evaluation. The primary metrics used for the evaluation of prediction sets are set size (average length of prediction sets; small value means high efficiency) and marginal coverage rate (fraction of testing examples for which prediction sets contain the ground-truth labels). These two metrics can be formally represented as :
$\operatorname{Size}=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}|\mathcal{C}(\bm{x}_{i})|$
$\operatorname{coverage\ rate}=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}\mathbf{1}(y_{i}\in\mathcal{C}(\bm{x}_{i}))$
Conditional coverage rate. In this work, we propose an alternative metric to the SSCV criterion named Each-Size Coverage Violation (ESCV) that can be utilized for any number of classes, as shown below:
$\operatorname{ESCV}(\mathcal{C},K)=\sup_{j}{\max({0,(1-\alpha)-\frac{\left|\left\{i:y_{i}\in\mathcal{C}\left(\bm{x}_{i},y_{i},u_{i}\right),i\in\mathcal{J}_{j}\right\}\right|}{\left|\mathcal{J}_{j}\right|}})}$
where $\mathcal{J}_{\mathrm{j}}=\{i:|\mathcal{C}\left(\bm{x}_{i},y_{i},u_{i}\right)|=j\}$ and $j\in\{1,\dots,K\}$. Specifically, ESCV measures the most significant violation of prediction sets with each size. This metric is practical because it only requires the set size, and is suitable for any classification problem, spanning from binary classes to large classes.
4.2 Results
Table 1: Results of average set sizes on different datasets. We evaluate the performance of SAPS, APS, and RAPS by calculating the average set size across multiple models. It is evident that SAPS consistently outperforms APS and RAPS in various classification tasks, such as ImageNet, CIFAR-100, and CIFAR-10, and different significance levels ($\alpha=0.1,0.05$). Bold numbers indicate optimal performance.
@@ 4.2 Results
SAPS generates smaller prediction sets. In Table 1, the performance of set sizes and coverage rates for various classification tasks are presented. We can observe that the coverage rate of all conformal prediction methods is close to the desired coverage $1-\alpha$. At different significance levels (i.e., $0.1$ and $0.5$), the prediction set size is consistently reduced by SAPS for ImageNet, CIFAR-100 and CIFAR-10, compared to APS and RAPS. For example, when evaluated on ImageNet, SAPS reduces the average set size from $20.95$ of APS to $2.98$. Moreover, as the scale of the classification task increases, the efficiency improvement achieved by SAPS becomes increasingly evident. Overall, the experiments show that our method has the desired coverage rate and a smaller set size than APS and RAPS. Due to space constraints, we only report the average results of multiple models on various classification tasks in Table 1, and detailed results for each model are available in Appendix D.
[…]
Figure 3: (a) ESCV with different models on ImageNet with $\alpha=0.1$. A good conformal prediction algorithm should keep the y-axis (e.g., ESCV) small. The results show that SAPS outperforms RAPS on most models. (b) Set size under different difficulties on VGG16. Small sets are required for easy examples, while hard ones require large sets. For example, SAPS generates smaller sets than RAPS on easy examples, but with difficulty improving, the size of SAPS will be larger than SAPS. (c) Set size on ImageNet-V2 at $\alpha=0.1$.
SAPS acquires lower conditional coverage violation. In Figure 3, we demonstrate that SAPS not only outperforms in efficiency but also boosts the conditional coverage rate,i.e., ESCV. Given that our study primarily focuses on improving the efficiency of prediction sets, the comparison of ESCV is limited to SAPS and RAPS. The results, shown in Figure 3, demonstrate that for most models, SAPS would get smaller ESCV than RAPS. For example, on CLIP, SAPS reduces the ESCV from $0.9$ to $0.37$. In addition, on ImageNet, we can observe that the ESCV of SAPS for different models is more stable than RAPS. Specifically, the ESCV of SAPS can keep a low value on most models, but in the case of RAPS, the maximum ESCV even gets $0.9$. The detailed results on CIFAR-10 and CIFAR-100 are provided in Appendix E.
SAPS exhibits higher adaptation. Adaptation indicates the ability to adjust the size of the prediction set based on the complexity or difficulty of individual examples. In other words, the prediction sets should be small for easy examples but large for hard ones. In this work, we employ the rank of the ground-truth labels in the sorted softmax probabilities to denote the difficulty. For instance, examples with serious difficulty are assigned high ranks for their ground-truth labels. In Figure 3, the results show that the set size of SAPS has higher adaptation. Specifically, compared with RAPS, SAPS produces smaller sets for accurate predictions but larger sets for hard examples on VGG16. More results of different models are reported in Appendix F. Overall, We show that SAPS can improve the adaptation of prediction sets while maintaining small set sizes.
Experiments on distribution shifts. We also verify the effectiveness of our method on the new distribution, which is different from the training data distribution. Specifically, We divide the test dataset of ImageNet-V2 (Recht et al., 2019), which exhibits a distribution shift compared to the ImageNet, equally into a calibration set containing 5000 images and a test set containing 5000 images. Then, the test models are only pre-trained on ImageNet and not be fine-tuned. As shown in Figure 3, the result shows that under $\alpha=0.1$, our method can also generate the smallest sets when the conformal calibration set and the test set come from a new distribution.
Table 2: Set size and ESCV for RAPS ($k_{r}=1$) and SAPS. We report the average value across various models with $\alpha=0.1$. The detailed results of each model are provided in the Appendix G. Bold numbers indicate optimal performance.
. Coverage Size $\downarrow$ ESCV $\downarrow$ Datasets RAPS($k_{r}=1$) SAPS RAPS($k_{r}=1$) SAPS RAPS($k_{r}=1$) SAPS ImageNet 0.900 0.900 3.24 2.98 0.631 0.396 CIFAR-100 0.899 0.899 2.79 2.67 0.390 0.302 CIFAR-10 0.900 0.898 1.62 1.63 0.138 0.089
@@
Figure 4: (a) Effect of the $\lambda$ on set size across various models. (b) Effect of the calibration dataset size on set size across various models. (c) Relationship between temperature and the set size of SAPS on ResNet152, where the horizon axis represents the log transformation of temperature $T$.
@@ 5 Discussion
Effect of $\lambda$ and calibration size. In SAPS, we choose an optimal $\lambda$ by performing a search over an arithmetic sequence (i.e., 0.02 to 0.6, spaced 0.03 apart) to minimize the set size on a validation set. In this work, the validation set constitutes 20% of the calibration set. Here, we provide an empirical analysis to show whether set size is sensitive to $\lambda$ and calibration size. To this end, we conduct two experiments on ImageNet to analyze the effects of $\lambda$ and the size of the calibration set.
@@
We present the results of four models in Figure 4. Indeed, Figure 3(a) illustrates that one can efficiently utilize grid search to find the optimal $\lambda$. Furthermore, as depicted in Figure 3(b), nearly all models maintain stable results when the number of calibration sets increases. Overall, the results demonstrate that the set size is not sensitive to variations in $\lambda$ and calibration size.
@@
SAPS vs. RAPS ($k_{r}=1$). While SAPS has demonstrated strong promise, it shares a similarity in the definition of non-conformity scores with RAPS ($k_{r}=1$), as shown below:
$S(\bm{x},y,u,\hat{\pi})=\sum_{i=1}^{o(y,\hat{\pi}(\bm{x}))-1}\hat{\pi}_{(i)}(\bm{x})+u*\hat{\pi}_{o(y,\hat{\pi}(\bm{x}))}(\bm{x})+\phi\cdot(o(y,\hat{\pi}(\bm{x}))-k_{r})^{+}.$
@@
Here, $\phi$ represents the weight of regularization and $(z)^{+}$ denotes the positive part of $z$. To this end, we conduct a comprehensive experiment with $\alpha=0.1$ on CIFAR-10, CIFAR-100, and ImageNet to compare SAPS and RAPS ($k_{r}=1$).
@@
As indicated in Table 2, SAPS outperforms RAPS ($k_{r}=1$) in large-scale classification scenarios, achieving smaller prediction sets and lower conditional coverage violations. In the small-scale classification task (i.e. CIFAR-10), SAPS produces a comparable set size with RAPS ($k_{r}=1$), and the ESCV of SAPS was more than 1.5 times as small as those from RAPS. Overall, employing a constant to substitute the noisy probabilities is an effective way to alleviate the negative implications of noisy probabilities further.
@@
Relation to temperature scaling. In the literature, temperature scaling calibrates softmax probabilities output by models by minimizing Expected Calibration Error (ECE), leading to a reliable maximum probability. As defined in Eq. 8, the remaining probability value used in the non-conformity scores is the maximum probability, a question arises: what is a relation between temperature scaling and the set size of SAPS? Here, we vary the value of temperature $T=~{}\{0.1,0.5,1,1.1,1.3,1.5,1.7,1.9,2,5,10,20\}$ in temperature scaling. We utilize SAPS to test the ResNet152 model calibrated by different temperatures on the ImageNet benchmark. The results indicate that there exists a consistency between the temperature value and the set size.
@@
As illustrated in Figure 3(c), the temperature with the lowest ECE can achieve the smallest prediction sets. Specifically, the optimal temperature on ECE and set size are the same, i.e., $1.3$. Moreover, as the ECE increases, the set size also increases. Indeed, temperature scaling can not change the permutation of the softmax probabilities but improves the confidence level of the maximum probability, resulting in the non-conformity scores of SAPS being more reliable. Overall, for SAPS, better confidence calibration can produce smaller prediction sets.

### v2

@@ 3.2 Method
In the above analysis, we demonstrate that removing the probability values in APS can largely decrease the size of prediction sets. Yet, the expected value of the set size (shown in Equation 8) will oscillate between $k-1$ and $k$, using APS without the probability value. This implies a shortcoming of the modified non-conformity score in adaptation to instance-wise uncertainty, leading to redundant prediction sets for easy examples.
@@
To alleviate this limitation, we propose a novel conformal prediction algorithm, named Sorted Adaptive Prediction Sets (SAPS). The key idea behind SAPS is to minimize the dependence of the non-conformity score on the probability values, while retaining the uncertainty information. In particular, we discard all the probability values except for the maximum softmax probability (MSP), which is usually used to measure the model confidence in the prediction (Hendrycks & Gimpel, 2016).
Formally, the non-conformity score of SAPS for a data pair $(\boldsymbol{x},y)$ can be calculated as
@@
$S_{saps}(\boldsymbol{x},y,u;\hat{\pi}):=\left\{\begin{array}[]{ll}u\cdot\hat{\pi}_{max}(\boldsymbol{x}),\qquad\text{if\quad}o(y,\hat{\pi}(\boldsymbol{x}))=1,\\ \hat{\pi}_{max}(\boldsymbol{x})+(o(y,\hat{\pi}(\boldsymbol{x}))-2+u)\cdot\lambda,\quad\text{else},\end{array}\right.$
@@
where $\lambda$ is a hyperparameter representing the weight of ranking information, $\hat{\pi}_{max}(\boldsymbol{x})$ denotes the maximum softmax probability and $u$ is a uniform random variable.
During inference, we calculate the non-conformity score for each label. In particular, the score is equivalent to that of APS for the label with the maximum softmax probability, while the scores of the resting labels are composed of the MSP and the ranking information from 2 to $o(y,\hat{\pi}(\boldsymbol{x}))$, i.e., $(o(y,\hat{\pi}(\boldsymbol{x}))-2+u)$. In this way, the new score of SAPS retains instance-wise uncertainty and mitigates the undesirable influence of tail probabilities.
With the new non-conformity score, the prediction sets of SAPS can be established according to Equation 6, following APS and RAPS. Consequently, we can conclude that SAPS also satisfies the marginal coverage guarantee, as stated in Theorem 2.1. Moreover, We provide an ablation experiment to investigate the relation between the prediction sets and the ranking weight $\lambda$ in Section 5.
Table 1: Performance comparison of various methods with different error rates $\alpha$. We conduct experiments on ImageNet and CIFAR-10/100 datasets. Bold numbers indicate optimal performance.
@@
4 Experiments
4.1 Experimental Setup
Datasets.
We consider three prominent datasets in our experiments: ImageNet (Deng et al., 2009), CIFAR-100 and CIFAR-10 (Krizhevsky et al., 2009), which are common benchmarks for conformal prediction. In particular, for ImageNet, its test dataset of 50,000 images is divided, allocating 30,000 images to the calibration set and 20,000 images to the test set. For both CIFAR-100 and CIFAR-10, the associated test dataset of 10,000 images is uniformly divided into two subsets: a calibration set and a test set, each comprising 5,000 images.
Models.
We employ twelve different classifiers, including nine standard classifiers, two transformer-based models, i.e., ViT (Dosovitskiy et al., 2020) and DeiT (Touvron et al., 2021), and a Vision-Language Model named CLIP (Radford et al., 2021). Aside from CLIP with zero-shot prediction capabilities, the remaining models are the models pre-trained on ImageNet. For CIFAR-10 and CIFAR-100, these models will be fine-tuned on the pre-trained models. Moreover, all classifiers are calibrated by the Temperature scaling procedure (Guo et al., 2017) before applying CP methods.
Conformal prediction algorithms.
We compare the proposed method against APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021). For methods that has hyper-parameters, we choose the hyper-parameter that achieves the smallest set size on a validation set, which is a subset of the calibration set. Specifically, we tune the regularization hyperparameter of RAPS in $\{0.001,0.01,0.1,0.15,\dots,0.5\}$ and hyperparameter $\lambda$ in $\{0.02,0.05,0.1,0.15,\dots,0.6\}$ for SAPS. All experiments are conducted with ten trials, and the median results are reported.
Additionally, we provide the details about the calibration process of conformal prediction algorithms, as shown:
  1. 1.
    Split: we split the full calibration set into a validation set and a calibration set ($20:80$ in this work);
  2. 2.
    Tuning: we use the validation set to choose proper hyperparameters of non-conformity score function;
  3. 3.
    Conformal Calibration: we use the calibration set to calculate the threshold $\tau$.
For CP methods with hyperparameters (i.e., RAPS and SAPS), we use the ”Split” and ”Tuning” steps to tune the hyperparameters. For APS, we use the full calibration set. Therefore, all methods have access to the same dataset for fair comparison.
@@ 4.1 Experimental Setup
Figure 2: (a) ESCV with different models on ImageNet with $\alpha=0.1$. A good conformal prediction algorithm should keep the y-axis (e.g., ESCV) small. The results show that SAPS outperforms RAPS on most cases. (b) Set size on ImageNet-V2 at $\alpha=0.1$. (c) Set size of various difficulties for multiple models on ImageNet. Small sets are required for easy examples, while hard ones require large sets.
Evaluation. The primary metrics used for the evaluation of prediction sets are set size (average length of prediction sets, small value means high efficiency) and marginal coverage rate (fraction of testing examples for which prediction sets contain the ground-truth labels). These two metrics can be formally represented as :
$\operatorname{Size}=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}|\mathcal{C}(\boldsymbol{x}_{i})|$
$\operatorname{Coverage\ rate}=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}\mathbf{1}(y_{i}\in\mathcal{C}(\boldsymbol{x}_{i}))$
Conditional coverage rate. In this work, we propose an alternative metric to the Size-stratified Coverage Violation (Angelopoulos et al., 2021) named Each-Size Coverage Violation (ESCV), which is given by:
$\operatorname{ESCV}(\mathcal{C},K)=\sup\limits_{j}{\max({0,1-\alpha-\frac{\left|\{i\in\mathcal{J}_{j}:y_{i}\in\mathcal{C}\left(\boldsymbol{x}_{i}\right)\}\right|}{\left|\mathcal{J}_{j}\right|}})}$
where $\mathcal{J}_{j}=\{i:|\mathcal{C}\left(\boldsymbol{x}_{i}\right)|=j\}$ and $j\in\{1,\dots,K\}$. Specifically, ESCV measures the most significant violation of prediction sets. This metric is practical because it only requires the set size, and is suitable for any classification problem, spanning from binary classes to large classes.
4.2 Results
SAPS generates smaller prediction sets. In Table 1, the performance of set sizes and coverage rates for various classification tasks are presented. We can observe that the coverage rate of all conformal prediction methods is close to the desired coverage $1-\alpha$. At different significance levels (i.e., $0.1$ and $0.05$), the prediction set size is consistently reduced by SAPS for ImageNet, CIFAR-100 and CIFAR-10, compared to APS and RAPS. For example, when evaluated on ImageNet, SAPS reduces the average set size from $20.95$ of APS to $2.98$. Moreover, as the scale of the classification task increases, the efficiency improvement achieved by SAPS becomes increasingly evident. Overall, the experiments show that our method has the desired coverage rate and a smaller set size than APS and RAPS. Due to space constraints, we only report the average results of multiple models on various classification tasks in Table 1, and detailed results for each model are available in Appendix C.
Experiments on distribution shifts. We also verify the effectiveness of our method on the new distribution, which is different from the training data distribution. Specifically, we divide the test dataset of ImageNet-V2 (Recht et al., 2019), which exhibits a distribution shift compared to the ImageNet, equally into a calibration set containing 5000 images and a test set containing 5000 images. Then, the test models are only pre-trained on ImageNet and not be fine-tuned. As shown in Figure 2(b), the result shows that under $\alpha=0.1$, our method can also generate the smallest sets when the conformal calibration set and the test set come from a new distribution.
SAPS acquires lower conditional coverage violation. In Figure 2(a), we demonstrate that SAPS not only outperforms in efficiency but also boosts the conditional coverage rate, i.e., ESCV. Given that our study primarily focuses on improving the efficiency of prediction sets, the comparison of ESCV is exclusively conducted between SAPS and RAPS. The results, shown in Figure 2(a), demonstrate that on ImageNet, SAPS would get smaller ESCV than RAPS for most models. For example, on CLIP, SAPS reduces the ESCV from $0.9$ to $0.37$. In addition, on ImageNet, we can observe that the ESCV of SAPS for different models is more stable than RAPS. Specifically, the ESCV of SAPS can keep a low value on most models, but in the case of RAPS, the maximum ESCV even gets $0.9$. The detailed results on CIFAR-10 and CIFAR-100 are provided in Appendix D.
Moreover, to further confirm that SAPS achieves superior conditional coverage than RAPS, we report the results of Size-stratified Coverage Violation (SSCV). In this experiment, the coarse partitioning of the set sizes for SSCV is set as: 0-1, 2-3, 4-10, 11-100, and 101-1000, which follows the experimental setting of RAPS (Angelopoulos et al., 2021). The average SSCV of various models on three datasets are shown in Table 2, with the experimental setup for ImageNet-V2 being consistent with the description provided in the preceding paragraph. We can observe that SAPS outperforms RAPS in conditional coverage. For instance, on ImageNet, SAPS records an SSCV of $0.23$, which is $0.05$ lower than that of RAPS. Although SAPS has better conditional coverage than RAPS, note that SSCV calculates the worst conditional coverage over the human-defined enclosures of set sizes, where different partitions lead to inconsistent SSCV results. Consequently, avoiding the human factor may lead to fair comparison, such as ESCV.
Table 2: The average SSCV and ESCV of various models on three datasets. Bold numbers indicate optimal performance.
ImageNet ImageNet-V2 CIFAR100
Metrics RAPS/SAPS(ours)
[…]
@@ 5 Discussion
Table 3: Set size and ESCV for RAPS ($k_{r}=1$) and SAPS. We report the average value across various models with $\alpha=0.1$. The detailed results of each model are provided in the Appendix E. Bold numbers indicate optimal performance.
Coverage Size $\downarrow$ ESCV $\downarrow$
Datasets RAPS($k_{r}=1$) SAPS RAPS($k_{r}=1$) SAPS RAPS($k_{r}=1$) SAPS
ImageNet 0.900 0.900 3.24 2.98 0.631 0.396
CIFAR-100 0.899 0.899 2.79 2.67 0.390 0.302
CIFAR-10 0.900 0.898 1.62 1.63 0.138 0.089
((a))
((b))
((c))
((d))
Figure 3: (a) The average ground-truth label ranking under different maximum softmax probabilities. Higher $\hat{\pi}_{max}$ have smaller label ranking. (b) Effect of the $\lambda$ on set size across various models. The black markers (,,$\bigstar$,$\blacklozenge$,$\blacktriangle$,$\bullet$ , ◆ , ▲ , ∙) represent the results of APS without probability value. (c) Effect of the calibration dataset size on set size across various models. (d) Relationship between temperature and the set size of SAPS on ResNet152, where the horizon axis represents the log transformation of temperature $T$.
SAPS exhibits adaptation.
In the literature of conformal prediction, the size of prediction sets is expected to represent the inherent uncertainty of the classifier’s predictions. Specifically, prediction sets should be larger for hard examples than for easy ones (Angelopoulos & Bates, 2021; Angelopoulos et al., 2021). In this work, we employ the rank of the ground-truth labels in the sorted softmax probabilities to denote the difficulty. For instance, examples with serious difficulty are assigned high ranks for their ground-truth labels. For simplicity, we choose a relatively coarse partitioning of image difficulty: 1, 2-3, 4-10, 11-100, and 101-1000. Then, we conduct an experiment on ImageNet to analyze the set size of different difficulties for SAPS.
In Figure 2(c), a salient observation is that the set size rises with increasing difficulties for different models. For example, on the Inception model, SAPS enlarges the length of the prediction sets from approximately $3.5$ for the simplest examples to nearly $9$ for the hardest examples. Moreover, we notice that the classifier with a higher accuracy rate can have a smaller set size under different levels of difficulty compared to the classifier with a lower accuracy rate. For example, the ResNeXt101 model has a smaller set size than DenseNet161 across different difficulties.
Table 4: Set size about retained top-$k$ probability values on the ImageNet. We report the average set size across various models with $\alpha=0.1$. Bold numbers indicate optimal performance.
$k$
Model 1 2 3 4 5
ResNeXt101 1.80 1.94 2.03 2.23 2.42
DenseNet161 2.12 2.16 2.27 2.46 2.65
ViT 1.60 1.67 1.76 1.86 1.97
Inception 4.60 4.50 4.55 4.77 4.95
Retained probability values.
To mitigate the redundancy of prediction sets for easy examples, we retain MSP in the softmax probabilities, as shown in Equation 10. In this manner, the SAPS method can not only produce efficient prediction sets but also communicate instance-wise uncertainty, i.e., adaptation. To further understand why retaining MSP in SAPS can decrease the set size of easy examples, we present the ground-truth label ranking of different MSP for three different models on ImageNet in Figure 3(a). We can observe that MSP is negatively related to ground-truth label ranking. Therefore, the ground-truth labels of easy examples (high MSP values) are generally highly ranked, which are expected to give compact prediction sets. In other words, the maximum softmax probability of a sample contains the information of its difficulty, allowing SAPS to produce adaptive prediction sets.
Here, we further discuss the effect of other softmax probability values on set size. Specifically, we retain the first $k$ maximum softmax probability values, where $k\in\{1,2,3,4,5\}$ We conduct experiments on the ImageNet with four classifiers of different architectures. In Figure 3(a), we show that retaining MSP achieves the best performance on most models. Additionally, more unreliable probability values involved in the score function will lead to larger prediction sets. Overall, retaining MSP in SAPS is the key to decreasing set size and preserving uncertainty.
To explore the sufficiency of MSP for conformal prediction, we provide an ablation study with some variants of SAPS by maintaining more probability values. Specifically, we compare five variants that preserve the top $k\in\{1,2,3,4,5\}$ values of softmax probabilities, respectively. The results of ImageNet using the four classifiers are reported in Table 4 of the manuscript. The results show that the vanilla SAPS with only MSP can achieve the best performance in most cases. For instance, with the ResNeXt101 model, only keeping the maximum softmax probability boosts the set size of $k=5$ from $2.42$ to $1.80$, a $0.62$ of direct improvement. In addition, retaining more tail-end probability values in the score function leads to larger prediction sets. Overall, retaining MSP is sufficient for SAPS.
Effect of $\lambda$ and the calibration size.
@@
In SAPS, we choose an optimal $\lambda$ by performing a search over a sequence to minimize the set size on a validation set. In this work, the validation set constitutes 20% of the calibration set. Here, we provide an experimental procedure to show whether set size is sensitive to $\lambda$ and calibration size. To this end, we conduct two experiments on ImageNet to analyze the effects of $\lambda$ and the size of the calibration set.
@@
We present the results of four models in Figure 3. Indeed, Figure 3(b) illustrates that one can efficiently utilize grid search to find the optimal $\lambda$. Furthermore, as depicted in Figure 3(c), nearly all models maintain stable results when the number of calibration sets increases. Overall, the results demonstrate that the set size is not sensitive to variations in $\lambda$ and the calibration size.
SAPS vs. RAPS ($k_{r}=1$).
@@
While SAPS has demonstrated strong promise, it shares a similarity in the definition of non-conformity scores with RAPS ($k_{r}=1$), as shown below:
$S_{raps}(\boldsymbol{x},y,u,\hat{\pi})$ $=\sum_{i=1}^{o(y,\hat{\pi}(\boldsymbol{x}))-1}\hat{\pi}_{(i)}(\boldsymbol{x})+u*\hat{\pi}_{o(y,\hat{\pi}(\boldsymbol{x}))}(\boldsymbol{x})$
$+\phi\cdot(o(y,\hat{\pi}(\boldsymbol{x}))-k_{r})^{+}.$
@@
Here, $\phi$ represents the weight of regularization and $(z)^{+}$ denotes the positive part of $z$. To this end, we conduct experiments on CIFAR-10, CIFAR-100, and ImageNet to compare SAPS and RAPS ($k_{r}=1$), with $\alpha=0.1$.
@@
As indicated in Table 3, SAPS outperforms RAPS ($k_{r}=1$) in large-scale classification scenarios, achieving smaller prediction sets and lower conditional coverage violations. In the small-scale classification task (i.e., CIFAR-10), SAPS produces a comparable set size with RAPS ($k_{r}=1$), and the ESCV of SAPS was more than 1.5 times as small as those from RAPS. Overall, employing a constant to substitute the noisy probabilities is an effective way to alleviate the negative implications of noisy probabilities further.
Relation to temperature scaling.
@@
In the literature, temperature scaling calibrates softmax probabilities output of models by minimizing Expected Calibration Error (ECE), leading to a reliable maximum probability (Guo et al., 2017). Moreover, as defined in Equation 10, the remaining probability value used in the non-conformity score function is the maximum probability. Thus, a question arises: what is a relation between temperature scaling and the set size of SAPS? Here, we vary the value of temperature $T=~{}\{0.1,0.5,1,1.1,1.3,1.5,1.7,1.9,2,5,10,20\}$ in temperature scaling. We utilize SAPS to test the ResNet152 model calibrated by different temperatures on the ImageNet benchmark. The results indicate that there exists a consistency between the temperature value and the set size.
@@
As illustrated in Figure 3(d), the optimal temperature on ECE, i.e., $1.3$coincides with the temperature that yields the smallest set size. Moreover, we can observe that an increase in ECE results in an expanded set size. Indeed, temperature scaling can not change the permutation of the softmax probabilities but improves the confidence level of the maximum probability, resulting in the non-conformity scores of SAPS being more reliable. Overall, for SAPS, confidence calibration is important to generate compact prediction sets.